OpenStudy (anonymous):

Find the derivative of f(x)=7 by the LIMIT PROCESS

OpenStudy (anonymous):

you need a function with a varible x otherwise it is not derivable

hartnn (hartnn):

lol it is differentiable

OpenStudy (anonymous):

\[f \prime(x)= (f(h+x)-f(x)) \div h\] hm, so how do I use this formula????

hartnn (hartnn):

what is f(x+h) can u find it ?

OpenStudy (anonymous):

.. not understadning your question.

OpenStudy (anonymous):

oops, I forgot to write the limit as h aprouched 0

hartnn (hartnn):

f(x) = 7 to find f(x+h) replace all 'x' in f(x) by 'x+h'

OpenStudy (anonymous):

like, 7+h?

hartnn (hartnn):

replace the 'x' do u find any 'x' in f(x) ?

OpenStudy (anonymous):

Do u mean in the 7 part? no

hartnn (hartnn):

so there is no x in f(x) so f(x+h) = 7 also did u get this ?

OpenStudy (anonymous):

oh, yes I did. But what is h? do I leave h as h???

OpenStudy (anonymous):

So, the slope is 0. does that mean its not differentiable?

hartnn (hartnn):

leave h as h only. so u get f(x+h) - f(x) =7-7 = 0 and slope =0 does not mean that its not differentiable

OpenStudy (anonymous):

I'm a little bit lost with this part. Im trying to find the limit as h aprouches 0. the entire equation is over h. so it comes down to Limit as h approuches 0 of (0/0)=0 does that mean the derivative is 0?

hartnn (hartnn):

since h-> 0 h is very very near to 0, but not equal to 0 like 0.00000000000001 or -0.99999999999999999 so 0/ h will be 0 hence derivative will be 0

OpenStudy (anonymous):

Awesome. The only thing I am stilff confused about is the f(x+h) part. What would I put there I was finding the derivative of something like f(x)=3x+2

hartnn (hartnn):

replace all x by 'x+h' so f(x+h) = 3(x+h) +2 ok ?

OpenStudy (anonymous):

oh! I get it! Thank you sooo sooo sooo much! I have big smile on my face. THe best way I can express that is throught these ---> :D :D :D THANK YOU!

hartnn (hartnn):

welcome ^_^ i love when people smile because of me <3

OpenStudy (anonymous):

Keep doing what ur doing! You rock!

hartnn (hartnn):

i will, thanks :)